OpenStudy (anonymous):

Write an equation in standard form for a line with a slope of –1 passing through C(2, 1).

OpenStudy (anonymous):

Use Point slope form: y-y1=m(x-x1) y-1=-1(x-2) y-1=-1x+2 y=-1x+3

OpenStudy (anonymous):

so in standard for would it be -1x+y=3 or 1x+y=3?

OpenStudy (anonymous):

Sorry, Ax+By=C, so its going to be 1x+y=3 (because from y=-1x+3 we are going to add the 1x to each side)

OpenStudy (anonymous):

or x+y=3, same thing